### v3

# EUvsDisinfo: a Dataset for Multilingual Detection of Pro-Kremlin Disinformation in News Articles

Abstract.
This work introduces EUvsDisinfo, a multilingual dataset of disinformation articles originating from pro-Kremlin outlets, along with trustworthy articles from credible / less biased sources. It is sourced directly from the debunk articles written by experts leading the EUvsDisinfo project. Our dataset is the largest to-date resource in terms of the overall number of articles and distinct languages. It also provides the largest topical and temporal coverage. Using this dataset, we investigate the dissemination of pro-Kremlin disinformation across different languages, uncovering language-specific patterns targeting certain disinformation topics. We further analyse the evolution of topic distribution over an eight-year period, noting a significant surge in disinformation content before the full-scale invasion of Ukraine in 2022. Lastly, we demonstrate the dataset’s applicability in training models to effectively distinguish between disinformation and trustworthy content in multilingual settings.
1. Introduction
Information warfare is a special case of disinformation that aims at manipulating public opinion to achieve military or political objectives (Thornton, 2015; Taylor, 2013; Lanoszka, 2016). Specifically, pro-Kremlin information warfare has been a topic of study across different disciplines (Abrams, 2016; Renz, 2016; Pamment and Bjola, 2016; Linvill et al., 2019). Since the annexation of Crimea, the spread of pro-Kremlin disinformation in Europe has intensified, encompassing many languages and even influencing search engines (Ižak, 2019; Lemke and Habegger, 2022; Jarynowski et al., 2023; Williams and Carley, 2023).
Dataset Size Labelling # Languages Multilinguality Topic(s) Period TALLIP-FakeNews (De et al., 2021) Manual Translation-based COVID-19 4 months MM-Covid (Li et al., 2020) Fact-checkers + Manual Inherent COVID-19 6 months FakeCovid (Kishore Shahi and Nandini, 2020) Fact-checkers Inherent COVID-19 4 months Multiverse (Dementieva et al., 2023) Source-based Semi-translation-based Celebrities, Science, Politics, Culture, World Unknown EUvsDisinfo (ours) Fact-checkers Inherent 508 topics originating from Pro-Kremlin disinformation 8.5 years
Combating pro-Kremlin disinformation articles presents significant challenges due to its dissemination across various languages, during an extensive period of time, and covering a wide range of different topics. Current disinformation datasets are not suitable for training machine learning models aimed to address this task since they lack the combination of such characteristics. Most datasets focus on analysing short claims (Thorne et al., 2018; Wang, 2017) or social media posts (Potthast et al., 2018; Zubiaga et al., 2016), whilst the majority of the existing article-level disinformation corpora consist of monolingual data (Hossain et al., 2020; Shu et al., 2020; Karadzhov et al., 2017; Hardalov et al., 2016; Wang et al., 2020; Golbeck et al., 2018; Pérez-Rosas et al., 2018; Shu et al., 2020; Shahi et al., 2021). A significantly smaller amount of multilingual datasets is available (De et al., 2021; Dementieva et al., 2023; Li et al., 2020; Kishore Shahi and Nandini, 2020), and they contain the following limitations: (i) representing parallel translated data (De et al., 2021; Dementieva et al., 2023), (ii) covering short periods of time ( 1 year) (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020), (iii) containing only few articles (around ) (De et al., 2021), and (iv) focusing on narrow topics (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020). To the best of our knowledge, the work by Solopova et al. (2023) is the only dataset addressing the theme of pro-Kremlin information warfare. Nevertheless, their dataset is targeted towards the detection of a pro-Western vs pro-Kremlin stance rather than disinformation.
In this work, we leverage the journalistic investigation carried out by specialists to generate a large and diverse multilingual dataset containing disinformation articles from pro-Kremlin outlets and trustworthy counterparts from reliable / less biased sources. To produce such a dataset, we use the debunk articles written by EUvsDisinfo111https://euvsdisinfo.eu, who since 2015 has been providing an EU-wide response to information warfare by debunking disinformation narratives across EU-member states. Our dataset is the largest ( articles) to-date, most topically diverse ( topics manually assigned by EUvsDisinfo), spans over the longest period of time ( years), and is the most diverse in terms of languages ( languages) compared to other article-level multilingual disinformation detection datasets (see Table 1 for an overview of related datasets in comparison to ours). Our key contributions are: (i) The novel multilingual disinformation dataset222Dataset: https://doi.org/10.5281/zenodo.10514307, becoming the largest and most diverse article-level set in terms of language, topics, and time periods. (ii) The analysis of the prominence of pro-Kremlin disinformation topics across different languages and time periods. (iii) The evaluation of several baseline models using our novel dataset for the task of multilingual binary disinformation detection. Our dataset and code are made publicly available, along with documentation, as supplementary material to facilitate reproducibility333Supplementary material: https://github.com/JAugusto97/euvsdisinfo.
2. Methodology
Our dataset leverages the structured information of the debunking articles published by EUvsDisinfo (see Figure 1 for an example). It consists of two classes of articles: disinformation and trustworthy. Disinformation articles are acquired directly from the links mentioned on the left hand-side of the EUvsDisinfo debunk page. Trustworthy articles are derived from the links to reliable sources within the response section of the article, in which EUvsDisinfo directly debunks the false narrative. Such URLs can lead to news articles, official documents, statements, books, encyclopedia articles, social media posts, and fact-checking articles, among others.
To verify that EUvsDisinfo does not reference disinformation articles in the response section, we manually inspected a randomly sampled set of 30 debunk articles, and annotated all the URLs mentioned within their response section. In total, 350 URLs were labelled as either trustworthy or potential disinformation. To do so, we considered the context in which the URL is referenced in the response section. For example, if the URL is referenced after a sentence such as ”see similar cases”, it is marked as potential disinformation (for reference, see the last paragraph in the response section of Figure 1). We identified that all of the URLs marked as potential disinformation actually correspond to other debunk articles by EUvsDisinfo or other fact-checking agencies such as Bellingcat and StopFake, instead of directly links to disinformation.
2.1. Data Collection
Given the wide variety of websites mentioned in EUvsDisinfo debunks, with varying HTML structure, extracting textual content is not trivial. For this task, we employ the Diffbot API444https://www.diffbot.com, which is a proprietary tool that uses machine learning to extract the content of a given web page. Although Diffbot is a closed-source service, it is free of charge for academic purposes. We retrieve articles that are no longer available on the web, by using the digital library Wayback Machine555https://web.archive.org. Also, the language of each trustworthy article is inferred using Polyglot666https://github.com/aboSamoor/polyglot, as EUvsDisinfo only specifies the languages of the disinformation articles.
Upon collecting the contents of all the URLs mentioned on EUvsDisinfo debunks, resulting in articles, we apply three filtering strategies to ensure consistency in the dataset. First, we remove URLs whose domain is not of a news outlet type. This is done to ensure consistency in the theme of disinformation detection of news articles published by news outlets. To do so, we first manually inspect the most frequent domains of the URLs, and label them as one of the following: “News Outlet”, “Organisation”, “Social Media”, “Fact-checker”, and “Other”. Next, we remove instances referring to error messages, log-in prompts, and paywalls through lexicon-bases rules, and through removing articles with less than characters, which is significantly shorter than the average article ( characters). Lastly, we remove URLs cited within sentences containing n-grams referring to other fact-checking articles (e.g. ”See earlier disinformation cases”). The full list of n-grams can be found in the supplementary materials\@footnotemark.
2.2. Dataset Overview
The full dataset has news articles with () and () marked as disinformation and trustworthy, respectively. It contains articles by different publishers on unique topics, spanning across years, between 06/01/2015 and 01/08/2023. On average, the articles are characters in length and cover different topics. The dataset contains different languages, of which are not present on existing multilingual article-level disinformation datasets presented in Table 1. Please refer to the supplementary material for a detailed breakdown of all languages. Out of all publishers in the dataset, and are associated with disinformation and trustworthy articles, respectively. The top five publishers of disinformation articles are known pro-Kremlin outlets: Sputnik ( ), RT (), RIA Novosti (), Tsargrad (), and Ukraina.ru (). For trustworthy articles, the top five publishers include Reuters (), BBC (), The Guardian (), Deutsche Welle (), and Radio Free Europe ().
2.3. Analysis of Topics
Figure 2 shows the distribution of the eight most frequent topics across the whole dataset for the five most common languages. Notably, the topic of “War in Ukraine” is the most prevalent across all five languages. However, certain topics appear more frequently in specific languages. The topic “Invasion of Ukraine” is over twice more common in Spanish than in other languages. “Conspiracy theory” and “Crimea” are frequently discussed in four out of the top five languages, except Spanish, where they account for only and of articles respectively, compared to an average of and in other languages. A similar discrepancy is noted for ”Invasion of Ukraine” in German, appearing in just of articles, while averaging in Russian, Spanish, Arabic, and English. These disparities suggest a language-specific targeted dissemination of disinformation contingent on the topics.
To examine the emergence of related themes over time, we group similar high-frequency topics into four overarching themes: “COVID-19”, “West”, “Russia”, and “Ukraine”. These themes are manually defined by merging similar topics from the top 50 most frequent in the dataset. Figure 3 shows the temporal distribution of disinformation topics under each theme. From the first quarter of 2021 onwards, there is a noticeable increase in disinformation articles across all themes, coinciding with Russia’s escalating military presence along the Ukrainian border. Similar surge in the number of “COVID-19”-related topics can be observed in the months following the start of the pandemic. After the full-scale invasion of Ukraine in the first quarter of 2022, disinformation articles on all topics decline significantly, except for ”War in Ukraine” and ”Invasion of Ukraine”, which become the dominant disinformation topics.
3. Experimental Setup
We employ the EUvsDisinfo dataset to perform binary disinformation detection (i.e. classify news articles as either disinformation or trustworthy). While the full dataset contains articles in unique languages, several languages are severely underrepresented and/or highly skewed towards one of the classes. Therefore in order to draw reliable conclusions, we discard underrepresented languages (fewer than articles) and those in which more than 95% of the articles belong to only one of the two classes. The resulting dataset used in the experiments comprises articles distributed across languages (% of disinformation articles is also shown)\@footnotemark: English (, ), Russian (, ), German (, ), French (, ), Spanish (, ), Georgian (, ), Czech (, ), Polish (, ), Italian (, ), Lithuanian (, ), Romanian (, ), Slovak (, ), Serbian (, ), Finnish (, ).
We experiment with four models that have been used in prior work on article-level disinformation classification. We train commonly adopted baselines from previous work (Shu et al., 2020; Golbeck et al., 2018; Li et al., 2020; Solopova et al., 2023): Multinomial Naive Bayes (MNB), and Support Vector Machine (SVM), both using bag-of-words to encode the textual features. We also train transformer-based models, mBERT and XLM-RoBERTa, as three of the relevant multilingual datasets discussed in Section 1 use mBERT as a baseline (De et al., 2021; Solopova et al., 2023; Kishore Shahi and Nandini, 2020), and XLM-R is used by Li et al. (2020). We conduct a hyperparameter search for each model by: (i) splitting the dataset into train and development sets with and of the data, respectively. The highest scoring hyperparameter configuration with respect to the development set is used throughout the experiments. (ii) discarding the development set, and splitting the remaining of the dataset into train and test sets using a 10-fold cross-validation strategy. The folds are stratified by both language and class. Model performance for each language is measured using the score to account for the skewed distribution of classes. To obtain a unified metric that encapsulates the overall system performance across all languages, we further average the scores per language to obtain the score. Further details such as hyperparameters and hardware configurations can be found in the supplementary material\@footnotemark.
4. Results
| Language | MNB | SVM | mBERT | XLM-R |
| EN | ± | ± | ± | ± |
| RU | ± | ± | ± | ± |
| DE | ± | ± | ± | ± |
| FR | ± | ± | ± | ± |
| ES | ± | ± | ± | ± |
| KA | ± | ± | ± | ± |
| CZ | ± | ± | ± | ± |
| PO | ± | ± | ± | ± |
| IT | ± | ± | ± | ± |
| LT | ± | ± | ± | ± |
| RO | ± | ± | ± | ± |
| SK | ± | ± | ± | ± |
| SR | ± | ± | ± | ± |
| FI | ± | ± | ± | ± |
| AVG |
Table 2 shows the classification results for the proposed baselines. mBERT achieves the highest average score (), and the highest per-language score in out of the languages: English, Russian, German, Spanish, Polish, Italian, Lithuanian, and Serbian. Next, XLM-R achieves an score of , which is lower than that of mBERT, despite XLM-R having roughly times more parameters. The standard deviations for mBERT and XLM-R show that their scores largely overlap for most languages. Nonetheless, XLM-R achieves the highest scores for Czech, Romanian, Slovak, and Finnish. The SVM baseline achieves an score of , a decrease of compared to mBERT. Surprisingly, the SVM baseline achieves the highest scores for two languages: French and Georgian. Lastly, the MNB model scores the lowest with an score of , a significant decrease compared to mBERT.
5. Conclusions and Future Work
This paper introduced EUvsDisinfo, a large, linguistically, temporally, and topically diverse dataset of disinformation and trustworthy articles originating from pro-Kremlin and reliable / less biased outlets, respectively. Using the dataset, we found evidences of language-specific targeting of specific topics, and revealed a surge in disinformation content related to the war in Ukraine right before its full-scale invasion in 2022. Lastly, we proposed classification baselines using our dataset for the task of binary disinformation detection in a multilingual setting. In future work, we plan to leverage the structure of our dataset to explore evidence-aware fact-checking approaches by linking disinformation and trustworthy articles referring to the same narrative.
Ethical Statement
The dataset is seeded from the publicly available debunks published by EUvsDisinfo. The authors of the manuscript are not part of the EUvsDisinfo organisation. The dataset content originates from news articles, thus it does not contain personal data. The dataset specifically targets pro-Kremlin disinformation; while pro-Western disinformation is not within the scope of this work. We recognise that the dataset is susceptible to misuse for malicious purposes (e.g., used by originators of disinformation to improve their techniques), and we strongly urge researchers to use it in accordance with best practice ethics protocols. Our dataset is compliant with FAIR principles (Wilkinson et al., 2016). It is made fully available with a unique digital object identifier, in a CSV format that can be processed by most widely used tools, and is released under an Apache 2.0 license. Since the textual content of news articles may be copyrighted, we do not include them in the dataset, but instead provide a software (https://doi.org/10.5281/zenodo.10492913) to allow researchers to collect the content themselves.
Acknowledgements.
This work has been co-funded by the UK’s innovation agency (Innovate UK) grant 10039055 (approved under the Horizon Europe Programme as vera.ai, EU grant agreement 101070093) under action number 2020-EU-IA-0282.References
- (1)
- Abrams (2016) Steve Abrams. 2016. Beyond propaganda: Soviet active measures in Putin’s Russia. Connections 15, 1 (2016), 5–31.
- De et al. (2021) Arkadipta De, Dibyanayan Bandyopadhyay, Baban Gain, and Asif Ekbal. 2021. A Transformer-Based Approach to Multilingual Fake News Detection in Low-Resource Languages. ACM Trans. Asian Low-Resour. Lang. Inf. Process. 21, 1, Article 9 (nov 2021), 20 pages. https://doi.org/10.1145/3472619
- Dementieva et al. (2023) Daryna Dementieva, Mikhail Kuimov, and Alexander Panchenko. 2023. Multiverse: Multilingual Evidence for Fake News Detection. Journal of Imaging 9, 4 (2023), 77.
- Golbeck et al. (2018) Jennifer Golbeck, Matthew Mauriello, Brooke Auxier, Keval H Bhanushali, Christopher Bonk, Mohamed Amine Bouzaghrane, Cody Buntain, Riya Chanduka, Paul Cheakalos, Jennine B Everett, et al. 2018. Fake news vs satire: A dataset and analysis. In Proceedings of the 10th ACM Conference on Web Science. 17–21.
- Hardalov et al. (2016) Momchil Hardalov, Ivan Koychev, and Preslav Nakov. 2016. In Search of Credible News. In Artificial Intelligence: Methodology, Systems, and Applications, Christo Dichev and Gennady Agre (Eds.). Springer International Publishing, Cham, 172–180.
- Hossain et al. (2020) Md Zobaer Hossain, Md Ashraful Rahman, Md Saiful Islam, and Sudipta Kar. 2020. BanFakeNews: A Dataset for Detecting Fake News in Bangla. In Proceedings of the Twelfth Language Resources and Evaluation Conference, Nicoletta Calzolari, Frédéric Béchet, Philippe Blache, Khalid Choukri, Christopher Cieri, Thierry Declerck, Sara Goggi, Hitoshi Isahara, Bente Maegaard, Joseph Mariani, Hélène Mazo, Asuncion Moreno, Jan Odijk, and Stelios Piperidis (Eds.). European Language Resources Association, Marseille, France, 2862–2871. https://aclanthology.org/2020.lrec-1.349
- Ižak (2019) Štefan Ižak. 2019. Using the topic of migration by pro-kremlin propaganda: case study of Slovakia. Journal of Comparative Politics 12, 1 (2019), 53–70.
- Jarynowski et al. (2023) Andrzej Jarynowski, Łukasz Krzowski, and Stanisław Maksymowicz. 2023. Biological mis (dis)-information in the Internet as a possible Kremlin warfare.
- Karadzhov et al. (2017) Georgi Karadzhov, Pepa Gencheva, Preslav Nakov, and Ivan Koychev. 2017. We Built a Fake News / Click Bait Filter: What Happened Next Will Blow Your Mind!. In Proceedings of the International Conference Recent Advances in Natural Language Processing, RANLP 2017, Ruslan Mitkov and Galia Angelova (Eds.). INCOMA Ltd., Varna, Bulgaria, 334–343. https://doi.org/10.26615/978-954-452-049-6_045
- Kishore Shahi and Nandini (2020) Gautam Kishore Shahi and Durgesh Nandini. 2020. FakeCovid – A Multilingual Cross-domain Fact Check News Dataset for COVID-19. arXiv e-prints, Article arXiv:2006.11343 (June 2020), arXiv:2006.11343 pages. https://doi.org/10.48550/arXiv.2006.11343 arXiv:2006.11343 [cs.CY]
- Lanoszka (2016) Alexander Lanoszka. 2016. Russian hybrid warfare and extended deterrence in eastern Europe. International affairs 92, 1 (2016), 175–195.
- Lemke and Habegger (2022) Tobias Lemke and Michael W Habegger. 2022. Foreign Interference and Social Media Networks: A Relational Approach to Studying Contemporary Russian Disinformation. Journal of Global Security Studies 7, 2 (04 2022), ogac004. https://doi.org/10.1093/jogss/ogac004 arXiv:https://academic.oup.com/jogss/article-pdf/7/2/ogac004/43510245/ogac004.pdf
- Li et al. (2020) Yichuan Li, Bohan Jiang, Kai Shu, and Huan Liu. 2020. Toward A Multilingual and Multimodal Data Repository for COVID-19 Disinformation. In 2020 IEEE International Conference on Big Data (Big Data). 4325–4330. https://doi.org/10.1109/BigData50022.2020.9378472
- Linvill et al. (2019) Darren L Linvill, Brandon C Boatwright, Will J Grant, and Patrick L Warren. 2019. “THE RUSSIANS ARE HACKING MY BRAIN!” investigating Russia’s internet research agency twitter tactics during the 2016 United States presidential campaign. Computers in Human Behavior 99 (2019), 292–300.
- Pamment and Bjola (2016) James Pamment and Corneliu Bjola. 2016. Digital containment: Revisiting containment strategy in the digital age. Global Affairs (2016).
- Pérez-Rosas et al. (2018) Verónica Pérez-Rosas, Bennett Kleinberg, Alexandra Lefevre, and Rada Mihalcea. 2018. Automatic Detection of Fake News. In Proceedings of the 27th International Conference on Computational Linguistics, Emily M. Bender, Leon Derczynski, and Pierre Isabelle (Eds.). Association for Computational Linguistics, Santa Fe, New Mexico, USA, 3391–3401. https://aclanthology.org/C18-1287
- Potthast et al. (2018) Martin Potthast, Johannes Kiesel, Kevin Reinartz, Janek Bevendorff, and Benno Stein. 2018. A Stylometric Inquiry into Hyperpartisan and Fake News. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Iryna Gurevych and Yusuke Miyao (Eds.). Association for Computational Linguistics, Melbourne, Australia, 231–240. https://doi.org/10.18653/v1/P18-1022
- Renz (2016) Bettina Renz. 2016. Russia and ‘hybrid warfare’. Contemporary Politics 22 (06 2016), 1–18. https://doi.org/10.1080/13569775.2016.1201316
- Shahi et al. (2021) Gautam Kishore Shahi, Julia Maria Struß, and Thomas Mandl. 2021. Overview of the CLEF-2021 CheckThat! lab task 3 on fake news detection. Working Notes of CLEF (2021).
- Shu et al. (2020) Kai Shu, Deepak Mahudeswaran, Suhang Wang, Dongwon Lee, and Huan Liu. 2020. Fakenewsnet: A data repository with news content, social context, and spatiotemporal information for studying fake news on social media. Big data 8, 3 (2020), 171–188.
- Solopova et al. (2023) Veronika Solopova, Christoph Benzmüller, and Tim Landgraf. 2023. The Evolution of Pro-Kremlin Propaganda From a Machine Learning and Linguistics Perspective. In Proceedings of the Second Ukrainian Natural Language Processing Workshop (UNLP), Mariana Romanyshyn (Ed.). Association for Computational Linguistics, Dubrovnik, Croatia, 40–48. https://doi.org/10.18653/v1/2023.unlp-1.5
- Taylor (2013) Philip M Taylor. 2013. Munitions of the mind: A history of propaganda from the ancient world to the present era. In Munitions of the Mind. Manchester University Press.
- Thorne et al. (2018) James Thorne, Andreas Vlachos, Christos Christodoulopoulos, and Arpit Mittal. 2018. FEVER: a Large-scale Dataset for Fact Extraction and VERification. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 809–819. https://doi.org/10.18653/v1/N18-1074
- Thornton (2015) Rod Thornton. 2015. The changing nature of modern warfare: responding to Russian information warfare. The RUSI Journal 160, 4 (2015), 40–48.
- Wang (2017) William Yang Wang. 2017. “Liar, Liar Pants on Fire”: A New Benchmark Dataset for Fake News Detection. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), Regina Barzilay and Min-Yen Kan (Eds.). Association for Computational Linguistics, Vancouver, Canada, 422–426. https://doi.org/10.18653/v1/P17-2067
- Wang et al. (2020) Yaqing Wang, Weifeng Yang, Fenglong Ma, Jin Xu, Bin Zhong, Qiang Deng, and Jing Gao. 2020. Weak Supervision for Fake News Detection via Reinforcement Learning. Proceedings of the AAAI Conference on Artificial Intelligence 34, 01 (Apr. 2020), 516–523. https://doi.org/10.1609/aaai.v34i01.5389
- Wilkinson et al. (2016) Mark D Wilkinson, Michel Dumontier, IJsbrand Jan Aalbersberg, Gabrielle Appleton, Myles Axton, Arie Baak, Niklas Blomberg, Jan-Willem Boiten, Luiz Bonino da Silva Santos, Philip E Bourne, et al. 2016. The FAIR Guiding Principles for scientific data management and stewardship. Scientific data 3, 1 (2016), 1–9.
- Williams and Carley (2023) Evan M Williams and Kathleen M Carley. 2023. Search engine manipulation to spread pro-Kremlin propaganda. Harvard Kennedy School Misinformation Review (2023).
- Zubiaga et al. (2016) Arkaitz Zubiaga, Maria Liakata, Rob Procter, Geraldine Wong Sak Hoi, and Peter Tolmie. 2016. Analysing how people orient to and spread rumours in social media by looking at conversational threads. PloS one 11, 3 (2016), e0150989.